OpenStudy (anonymous):

Explain why (-81)^1/2 is not defined?

OpenStudy (amistre64):

becasue you havent defined it?

OpenStudy (zepp):

^(1/2) is the same thing as square root.

OpenStudy (anonymous):

because x^1/2 means the sqrt of x sqrt of a negative number is not defined in the real numbers

OpenStudy (zepp):

Hence, \(-81^{\frac{1}{2}} = \sqrt{-81}\)

OpenStudy (zepp):

And since a square root of a negative number doesn't belong to the set of rational numbers, we say that \(-81^{\frac{1}{2}}\) is undefined.

OpenStudy (anonymous):

Thanks zepp.

OpenStudy (anonymous):

Thanks thomasj.